QUESTION 6

  1. ABC & Co expects a net income of K 80,000. It has 8% Debentures worth K 2,000,000. The equity capitalisation rate of the company is 10%.
  1. Calculate the value of the firm and overall cost of capital according to the Net Income approach. (Ignoring income tax)
  2. If Debenture debt is increased to K 3,000, 000 what shall be the value of firm and the overall cost of capital?                                                             (12 marks)

.(III) SLM Ltd expects a net operating income of K 2,000,000. It has 6% Debentures worth K 10,00,000. The overall capitalisation rate is 10%.

  1. Calculate the value of the firm and the equity capitalisation rate according to NOI approach.
  2. If the debenture debt is decreased to K 7,500,000 what will be effect on the value of the firm and the equity capitalisation rate ?                              (13 marks)

(Total: 25 marks)
Answer #1

Answer:-

a) i) Net Income =K 80,000

Interest on debentures = (K 2,000,000 × 8%) = K 16,000

Earning available to equity shareholder= net income - interest on debentures

Earning available to equity shareholder =K 80,000 - K 16,000

Earning available to equity shareholder =K 64,000

Market value of the equity (S) = earning available to equity ×(100/10)

Market value of the equity(S) =64,000 (100/10) = K 6,400,000

Market value of the debenture(D) = K 2,000,000

Value of the firm (S +D) = K 6,400,000 + K 2,000,000

Value of the firm = K 8,400,000

Overall cost of capital =(Net income/value of firm) × 100

Overall cost of capital = (K 80,000 /K 8,400,000)×100

Overall cost of capital = 9.52%

a) ii) If Debenture debt is increased to K 3,000, 00

Net Income =K 80,000

Interest on debentures = (K 3,000,000 × 8%) = K 24,000

Earning available to equity shareholder= net income - interest on debentures

Earning available to equity shareholder =K 80,000 - K 24,000

Earning available to equity shareholder =K 56,000

Market value of the equity (S) = earning available to equity ×(100/ equity capitalisation rate)

Market value of the equity(S) =56,000 (100/10) = K 5,600,000

Market value of the debenture(D) = K 3,000,000

Value of the firm (S +D) = K 5,600,000 + K 3,000,000

Value of the firm = K 8,600,000

Overall cost of capital =(Net income/value of firm) × 100

Overall cost of capital = (K 80,000 /K 8,600,000)×100

Overall cost of capital = 9.30%

As HomeworkLib rule, We only give answer for 1 questions only